# Einstein-Podolsky-Rosen correlations in spontaneous parametric down-conversion: Beyond the gaussian approximation

### v1

@@ V Discussion and Conclusion
From the results presented here, one can see that the EPR correlations $\Delta x\Delta k_{x}$ and $\Delta y\Delta k_{y}$ of the photons pairs generated by spontaneous parametric down-conversion are strongly affected by the pump beam angular spectrum in the direction normal to the principal plane (defined by the optic axis and the $z$ axis). By another side, such dependence is much smaller in the direction parallel to the principal plane. This effect is readily explained by the presence of the term $\operatorname{sinc}l_{t}Q_{x}$ in Eq. (II). That term, which depends on the birefringence, on the phase match angle, and on the crystal length (see Eq. (2)), acts as a spatial filter for the transfer of the angular spectrum from the pump beam to the two-photon state [17]. Because of this filtering effect, the product $\Delta x\Delta k_{x}$ behaves as if the laser beam were less focused. The two uncertainty products $\Delta x\Delta k_{x}$ and $\Delta y\Delta k_{y}$ tend to a unique minimum value as the pump beam gets more collimated, that is, $w_{0}\gg l_{t}$. In our case, $l_{t}=0.186\,\mathrm{mm}$.
In conclusion, we have presented: (a) Accurate analytic expressions for the coincidence detection probability amplitudes of photon pairs generated by spontaneous parametric down-conversion in both momentum and position spaces. Those expressions allow us to predict how the correlations in position and momentum depend on the system parameters like crystal length, crystal birefringence, pump beam focusing, pump beam waist location and detectors location. (b) Experimental data supporting our theoretical predictions, using Einstein-Podolsky-Rosen correlations as benchmarks, for 8 different pump beam configurations.

### v2

@@ V Discussion and Conclusion
From the results presented here, one can see that the EPR correlations $\Delta x\Delta k_{x}$ and $\Delta y\Delta k_{y}$ of the photons pairs generated by spontaneous parametric down-conversion are strongly affected by the pump beam angular spectrum in the direction normal to the principal plane (defined by the optic axis and the $z$ axis). Such dependence is much smaller in the direction parallel to the principal plane. This effect is readily explained by the presence of the term $\operatorname{sinc}l_{t}Q_{x}$ in Eq. (II). That term, which depends on the birefringence, the phase match angle, and the crystal length (see Eq. (3)), acts as a spatial filter for the transfer of the angular spectrum from the pump beam to the two-photon state [19]. Because of this filtering effect, the product $\Delta x\Delta k_{x}$ behaves like the laser beam was less focused. The two uncertainty products $\Delta x\Delta k_{x}$ and $\Delta y\Delta k_{y}$ tend to a unique minimum value as the pump beam gets more collimated, that is, $w_{0}\gg l_{t}$. In our case, $l_{t}=0.186\,\mathrm{mm}$.
In conclusion, we have presented, unlike any previous publication: (a) Accurate analytic expressions for the coincidence detection probability amplitudes of photon pairs generated by spontaneous parametric down-conversion in both momentum and position spaces on the entire plane normal to the pump beam. Those expressions allow us to predict how the correlations in position and momentum depend on the system parameters like crystal length, crystal birefringence, pump beam focusing, pump beam waist location, and detectors locations. (b) Experimental data supporting our theoretical predictions, using Einstein-Podolsky-Rosen correlations as benchmarks, for 8 different pump beam configurations.